data points in the temperature range 300 K to 3000 K. 1.14 Calculate the thermal radiation emitted by a sur- face with an emissivity of 0.8 and a temperature of 1300℃. Din 1.20 1.21Number 14
Answer #1

Thermal radiation is given as;

q=\epsilon \sigma T^4

here,

q = thermal\;\;radiation

\epsilon = emissivity

\sigma = Stefan-Boltzmann\;\;constant=5.67\times10^{-8}\;\;W/m^2K^4

T=temperature\;\;(in\;\;Kelvin)

_____________________________________________________________________

Given,

T = 1300^{\circ}C=(1300+273.15)\;\;K=1573.15\;\;K

\epsilon = 0.8

Hence,

q=\epsilon \sigma T^4

\Rightarrow q=(0.8) (5.67\times10^{-8}\;\;W/m^2K^4) (1573.15\;\;K)^4

\Rightarrow q=277813.6\;\;W/m^2

\Rightarrow q=2.778\times 10^5\;\;W/m^2